6) A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem, where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 3 feet and 4 feet.
Answer:
Show work in this space

Answer by funmath(2933)

The pythagorean theorem states that in a right triangle, (hypotenuse)^2=(adjacent leg)^2+(opposite leg)^2, otherwise stated:c^2=a^2+b^2
c^2=3^2+4^2
c^2=9+16
c^2=25
sqrt(c^2)=sqrt(25)
c=5 feet